The standard potentials (at pH = 7) are given for some important donors and
acceptors:
Reduced Species, Oxidized Product, Potential( V)
H2, H -041
H2O, 02, +0.822
H2 S, so4, -0.22
N2, No3, +0.75
Fe2+, Fe3+, +0.77
CH4 ,CO2, -0.24
Compute the overall standard potential (at pH 7) for transfer of electrons from
the follow pairs ofdonors and acceptors. Rank them from most energetic (most
positive) to least energetic: H2/Fe3+, H2S/02, CH4IN03", H2/02, Fe2+102,
and H2SIN03"
1
Expert's answer
2020-04-20T09:49:40-0400

For the reaction

E0r=EO(Ox1/Red1)-EO(Ox2/Red2)

-Eor=E0ox-E0red

E0r(H2/Fe3+)=+0,77-(-0,41)=1,18 V

E(H2S/O2)=+0,822-(-0,22)=-1,042V

E(CH4/NO3-)=+0,75-(-0,24)=0,99V

E(H2/O2)=0,822-(-0,41)=1,232V

E(Fe2+/O2)=0,822-0,77-=0.052V

E(H2S/NO3-)=+0.75-(-0.22)=0,97V